you are provided with 48cells of each 2 emf Nd internal resistance 4 ohms .what is maximum current flow through the circuit if total external resistance is 12 ohms

Grade:12

1 Answers

Kamiya
6 years ago
V= IR V/R = I N is number of cells V is emf of cellsR is esternal resistancer is internal resistanceNV/ R+Nr = I 48×2 / 12+48×4 =I2/3 ANS